Frequency of tuning forks are 320 Hz. And 325 Hz. If they are sounded together, the beat period is…….

(a) 5 second
(b) 6 second
(c) 0 second
(d) 0.2 second

Grade:Upto college level

1 Answers

NIRAV
6 years ago
beat frequency = f2-f1= 325-320= 5.
freuency is related inversly with time period.
time period = 1/(f2-f1) =1/5= 0.2 seconds